OpenStudy (anonymous):

find the input when the output is 17

OpenStudy (anonymous):

17

OpenStudy (anonymous):

the output of a function is 5 more than 2 times the input &then find the input when the output is 17

OpenStudy (anonymous):

6

OpenStudy (anonymous):

whut?

OpenStudy (whynot):

output=2(input)+5 17=2x+5 (17-5)/2=x x=12/2 x=6